If a savings account pays 3% simple annual interest, a deposit of $250 will earn in 2 years?
Answer by jim_thompson5910(35256)

Interest earned:


I = P*r*t


I = 250*0.03*2


I = 15


You will earn 15 dollars in interest.


So overall, your account will have 250+15 = 265 dollars in it. This is assuming no deposits or withdrawals are made within the 2 year period.